OpenStudy (anonymous):

f (x) = 3x/2x - 4 what is the domain function and write it in interval notation please help solve

OpenStudy (anonymous):

please post correctly if this isquestion then 3x/2x=3/2 then it is not the function of x

OpenStudy (anonymous):

I type it just like the instructor gave it to me let me see if I can copy and paste it

OpenStudy (anonymous):

identify the domain of this function and write the domain in interval notation f (x) = 3x/2x -4

OpenStudy (anonymous):

Hi are you there ?

OpenStudy (anonymous):

\[(-\infty,2) \cup(2,\infty)\]

OpenStudy (anonymous):

could you please provide more information how you solved the problem? so what is the domain function. I am sorry but this problem is new to me.

OpenStudy (anonymous):

2x-4=0 2x=4 x=2 this function is undefined,when x=2

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

the domain of rational function is the set of the values of the variables for which the denominator is not zero

OpenStudy (anonymous):

thank you so much